OpenStudy (anonymous):

How would you solve these equation systems with matrices? x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10

OpenStudy (anonymous):

i would get rid to the x first then solve for one of the other variables

OpenStudy (anonymous):

Ok, I think I'll buy a book and learn to do matrices on my own. ^^

OpenStudy (anonymous):

oh my bad, i didnt even read the who question

OpenStudy (anonymous):

i just thought it said solve the system

OpenStudy (anonymous):

Yeah, well I know how to do that. But I can't construct a matrix because I haven't done that before. It's interesting and new to me.

OpenStudy (anonymous):

this is what we have ( meaing the coefficient for each linear equation)

OpenStudy (anonymous):

You put the equations into one equation, witch is called a matrix right?

OpenStudy (anonymous):

1 4 1 -10 3 -3 6 -21 1 2 2 -10

OpenStudy (anonymous):

So if I put the numbers: 1 4 1 x -10 [3 -3 6 ] * [ y ] = [ -21 ] 1 2 2 z -10 Is that correct?

OpenStudy (anonymous):

Yes. But then you make a matrix like\[\left[\begin{matrix}1 & 4 & 1 & -10 \\ 3 & -3 & 6 & -21 \\ 1 & 2 & 2 & -10\end{matrix}\right]\]and then solve using Gaussian elimination.